# Accurate Small-Signal Modeling of Digitally Controlled Buck Converters with ADC-PWM Synchronization

### v3

@@ Abstract
Digital control has become increasingly widespread in modern power electronic converters. When acquiring feedback signals such as the inductor current, synchronizing the analog-to-digital converter (ADC) with the digital pulse-width modulator (DPWM) is commonly employed to accurately track their steady-state average. However, the small-signal implications of such synchronization have not been investigated. This paper presents an exact small-signal model for digitally controlled buck converters operating in forced continuous-conduction mode (FCCM) under constant-frequency current-mode control, explicitly accounting for DPWM-ADC synchronization. Using a sampled-data framework, the proposed model captures all sideband effects introduced by the sampling process, yielding precise predictions of both analog and digital loop gains, even at frequencies beyond the switching and sampling frequencies. Both asymmetrical and symmetrical carrier modulations are considered. Furthermore, the digital loop gain is derived in closed form using the modified z-transform, enabling low-complexity compensator design and stability assessment. Within this framework, the analog loop gain can be directly obtained from the digital loop gain, thereby eliminating the need for computationally intensive infinite series evaluations. The validity of the proposed model is confirmed through both simulation and experimental results.
@@ I Introduction
As digital control becomes increasingly popular in power converters, its advantages in cost, programmability, and integration over analog designs have made it the preferred choice. [1] In particular, it greatly simplifies the realization of sophisticated algorithms. Moreover, it enables over-the-air (OTA) firmware updates, allowing deployed products to gain new functionality without hardware modifications.
In common engineering practice, digital power converter design typically begins with power stage modeling using the classical state-space averaging (SSA) technique [2, 3]. An analog compensator is then designed in the continuous-time domain, discretized, and finally implemented in the digital controller. By averaging the piecewise linear state-spaces over one switching period, SSA eliminates the non-linearities. However, SSA neglects essential effects, such as computational delays, sampling, the presence of ripples, and sideband component coupling. [4] Consequently, SSA loses validity even if the perturbation frequency is far below the switching frequency. While SSA leverages well-established linear time-invariant (LTI) control theory for rapid design, it fails to provide an accurate prediction of the closed-loop stability, particularly when the crossover frequency approaches the sampling frequency, leading to a potential risk of instability.
@@
As a compromise between model accuracy and computational complexity, researchers have proposed simplified models that include only a limited number of sidebands. Two-frequency models proposed by [16] and [17] considered one sideband component other than the perturbed frequency $f_{p}$. By considering a pair of sideband components and the switching frequency component, [18] fixes the low-frequency phase error in two-frequency models. An extended frequency model that incorporates all sideband components to enhance accuracy under large ripple conditions was proposed in [19]. However, the analytical form of the transfer function was not provided in [19].
@@
The remainder of this paper is organized as follows. Section II presents the small-signal sampled-data modeling under asymmetrical carrier modulation with explicit consideration of DPWM-ADC synchronization. A closed-form expression of the digital loop gain is derived using the modified z-transform, and it is shown that the analog loop gain can be obtained from the digital loop gain, thereby avoiding complicated infinite-series evaluations. Although the derivation begins with trailing-edge modulation (TEM), the conclusions are also extended to leading-edge modulation (LEM). Section III develops the small-signal model for the case of a symmetrical carrier. Section IV compares the proposed model with the simulation results and existing models, demonstrating its superior accuracy. Section V provides experimental validation, confirming that the proposed model exhibits good agreements in all cases. Finally, Section VI concludes the paper.
@@ II Small-Signal Modeling of Digital Buck Converters with Asymmetrical Carriers
Fig. 2 illustrates a digitally controlled synchronous Buck converter employing current-mode control and operating in forced continuous conduction mode (FCCM). The digital counter generates an up-counting sawtooth carrier ($CNTR$) with an amplitude of $CNTR_{MAX}$. When $CNTR$ reaches zero, the counter updates the shadow register. The shadow register serves to suppress spurious PWM pulses during abrupt changes in the duty-cycle command. The DPWM output will be high if the shadow register output is higher than $CNTR$ and will be low elsewhere. This configuration implements a TEM DPWM, in which the PWM rising edge is fixed at the beginning of each cycle, while the falling edge varies and moves according to the duty-cycle command.
@@
Figure 3: Key discrete-domain waveforms of a TEM buck converter. Dense and sparse vertical dashed lines denote the sampling instants at steady state and under perturbation, respectively. The ADC samples at integer multiples of $T_{S}$.
@@
Since DPWM-ADC synchronization guaranties that sampling always occurs at the PWM on-interval center, as shown in Fig.4, in each cycle, the ADC sampling instant $t_{smp}$ is related to the on-time perturbation $\hat{t}_{on}$ of the previous cycle:
@@ III Small-Signal Modeling of Digital Buck Converters with Symmetrical Carriers
(25) provides the plant transfer function for digital compensator design. The open-loop transfer function in the digital domain can then be computed from (19), while the corresponding analog-domain loop gain can still be obtained directly using (20). In this case, $T_{0}(s)$ is modified as:
@@ IV Comparison Between Existing Small-Signal Models and the Proposed Analytical Model
Figure 9: Plant transfer function as perceived by the digital compensator (TEM + turn-off-centered sampling): the proposed model, Dragan’s model [30], and simulation.

### v4

@@ Abstract
Digital control has become increasingly widespread in modern power electronic converters. When acquiring feedback signals such as the inductor current, synchronizing the analog-to-digital converter (ADC) with the digital pulse-width modulator (DPWM) is commonly employed to accurately track their steady-state average. However, the small-signal implications of such synchronization have not been investigated. This paper presents an exact small-signal model for digitally controlled buck converters operating in forced continuous-conduction mode (FCCM) under constant-frequency current-mode control, explicitly accounting for ADC-DPWM synchronization. Using a sampled-data framework, the proposed model captures all sideband effects introduced by the sampling process, yielding precise predictions of both analog and digital loop gains, even at frequencies beyond the switching and sampling frequencies. Both asymmetrical and symmetrical carrier modulations are considered. Furthermore, the digital loop gain is derived in closed form using the modified z-transform, enabling low-complexity compensator design and stability assessment. Within this framework, the analog loop gain can be directly obtained from the digital loop gain, thereby eliminating the need for computationally intensive infinite series evaluations. The validity of the proposed model is confirmed through both simulation and experimental results.
@@ I Introduction
As digital control becomes increasingly popular in power converters, its advantages in cost, programmability, and integration compare to analog designs have made it the preferred choice. [1] In particular, it greatly simplifies the realization of sophisticated algorithms. Moreover, it enables over-the-air (OTA) firmware updates, allowing deployed products to gain new functionality without hardware modifications.
In common engineering practice, digital power converter design typically begins with power stage modeling using the classical state-space averaging (SSA) technique [2, 3]. An analog compensator is then designed in the continuous-time domain, discretized, and finally implemented in the digital controller. By averaging the piecewise linear state-spaces over one switching period, SSA eliminates the non-linearities. However, SSA neglects essential effects, such as computational delays, sampling, the presence of ripple, and sideband component coupling [4]. Consequently, SSA loses validity even if the perturbation frequency is far below the switching frequency. While SSA leverages well-established linear time-invariant (LTI) control theory for rapid design, it fails to provide an accurate prediction of the closed-loop stability, particularly when the crossover frequency approaches the sampling frequency, leading to a potential risk of instability.
@@
As a compromise between model accuracy and computational complexity, researchers have proposed simplified models that include only a limited number of sidebands. Two-frequency models proposed by [16] and [17] considered one sideband component other than the perturbed frequency $f_{p}$. By considering a pair of sideband components and the switching frequency component, [18] fixes the low-frequency phase error in two-frequency models. An extended frequency model that incorporates all sideband components to enhance accuracy under large ripple conditions was proposed in [19]; However, the analytical form of the transfer function was not provided.
@@
The remainder of this paper is organized as follows. Section II presents the small-signal sampled-data modeling under asymmetrical carrier modulation with explicit consideration of DPWM-ADC synchronization. A closed-form expression of the digital loop gain is derived using the modified z-transform, and it is shown that the analog loop gain can be obtained from the digital loop gain, thereby avoiding complicated infinite-series evaluations. Although the derivation begins with trailing-edge modulation (TEM), the conclusions are also extended to leading-edge modulation (LEM). Section III develops the small-signal model for the case of a symmetrical carrier. Section IV compares the proposed model with the simulation results and existing models, demonstrating its superior accuracy. Section V provides experimental validation, confirming that the proposed model exhibits good agreements in all cases. Section VI concludes the paper.
@@ II Small-Signal Modeling of Digital Buck Converters with Asymmetrical Carriers
Fig. 2 illustrates a digitally controlled synchronous Buck converter employing current-mode control and operating in forced continuous conduction mode (FCCM). The digital counter generates an up-counting sawtooth carrier $CNTR(t)$ with an amplitude of $CNTR_{MAX}$. When $CNTR$ reaches zero, the counter updates the shadow register. The shadow register serves to suppress spurious PWM pulses during abrupt changes in the duty-cycle command. The DPWM output will be high if the shadow register output is higher than $CNTR$ and will be low elsewhere. This configuration implements a TEM DPWM, in which the PWM rising edge is fixed at the beginning of each cycle, while the falling edge varies and moves according to the duty-cycle command.
@@
Figure 3: Discrete-domain waveforms of a TEM buck converter. Dense and sparse vertical dashed lines denote the sampling instants at steady state and under perturbation, respectively. The ADC samples at integer multiples of $T_{S}$.
@@
Since DPWM-ADC synchronization guaranties that sampling occurs exactly at the midpoint of the PWM on-interval, as shown in Fig.4, in each cycle, the ADC sampling instant $t_{smp}$ is related to the on-time perturbation $\hat{t}_{on}$ of the previous cycle:
@@ III Small-Signal Modeling of Digital Buck Converters with Symmetrical Carriers
Equation (25) provides the plant transfer function for digital compensator design. The open-loop transfer function in the digital domain can then be computed from (19), while the corresponding analog-domain loop gain can still be obtained directly using (20). In this case, $T_{0}(s)$ is modified as:
@@ IV Comparison Between Existing Small-Signal Models and the Proposed Analytical Model
Figure 9: Plant transfer function seen by the digital compensator (TEM + turn-off-centered sampling): the proposed model, Dragan’s model [30], and simulation.